# CASE – Condition-Aware Sentence Embeddings for Conditional Semantic Textual Similarity Measurement

### v3

@@ Abstract
We use the original C-STS training set and the re-annotated validation set in this paper. The dataset, especially the training set, includes annotation errors, such as invalid conditions and subjective human annotations. These disadvantages hurt the model’s ability to capture the semantics and perform the C-STS task. This further introduces ambiguity and inaccuracy to CASE. Manually re-annotating the training set, which contains 11342 instances, is costly and is deferred to future work.
Our evaluations cover only English, which is a morphologically limited language. To the best of our knowledge, C-STS datasets have not been annotated for languages other than English, which has limited all prior work on C-STS to conduct experiments using only English data. However, the sentence encoders we used in our experiments support multiple languages. Therefore, we consider it to be an important future research direction to create multilingual datasets for C-STS and evaluate the effectiveness of our proposed method in multilingual settings.
There is a large number of sentence encoders (over 240 models as at March 2025 evaluated on Massive Text Embedding Benchmark (MTEBMuennighoff et al. (2023) leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard). However, due to computational limitations, we had to select a subset covering the best performing (top ranked on MTEB at the time of writing) models for our evaluations. We will release our source code and the evaluation framework such that the research community can evaluate our proposed method with any sentence encoder for C-STS.
@@
We did not collect or annotate any datasets in this project. Instead, we use existing C-STS datasets annotated and made available by Deshpande et al. (2023) and Tu et al. (2024). To the best of our knowledge, no ethical issues have been raised regarding those datasets.
@@ Appendix A Models
Appendix B Isotropy for Embeddings
We first use the embedding-to-mean cosine similarity distribution to measure the isotropy of the embeddings. Given a set of embeddings $S=\{\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{n}\}$, we first compute the mean embedding vector $\bm{\mu}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}$. Then, for each embedding $\bm{x}_{i}\in S$, we compute its cosine similarity with the mean vector $\mu$, i.e., $\cos(\bm{x}_{i},\bm{\mu})=\frac{\bm{x}_{i}{}^{\top}\bm{\mu}}{\left|\left|\bm{x}_{i}\right|\right|\left|\left|\bm{\mu}\right|\right|}$. The distribution of these embedding-to-mean cosine similarities is then analysed to characterise the embedding space – a distribution sharply peaked near 1 indicates anisotropy, whereas a broader, more uniform distribution suggests a more isotropic geometry.
@@ Appendix B Isotropy for Embeddings
From Table 4, Figure 5, and Figure 6, we see that the embeddings after subtracting c have a lower mean cosine similarity to the mean vector and a higher standard deviation, indicating that they are more spread out in the embedding space. In contrast, the embeddings without subtracting c are more clustered around a central direction (higher mean, lower standard deviation), reflecting anisotropy, a tendency for vectors to concentrate in a narrow region. Therefore, embeddings after subtracting c tend to be more isotropic, indicating better distributional diversity.
@@
(c) SimCSE_base: $\operatorname{CONC}(c+s)$ c
@@
(e) SimCSE_large: $\operatorname{CONC}(c+s)$ c
@@
Model Embedding Type Mean Std
NV cond - c 0.407 0.084
cond 0.492 0.067
SFR cond - c 0.537 0.055
cond 0.708 0.021
GTE cond - c 0.489 0.067
cond 0.696 0.036
E5 cond - c 0.542 0.056
cond 0.897 0.010
SimCSE_base $\operatorname{CONC}(c+s)$ c 0.254 0.095
$\operatorname{CONC}(c+s)$ 0.347 0.088
SimCSE_large $\operatorname{CONC}(c+s)$ c 0.248 0.110
$\operatorname{CONC}(c+s)$ 0.333 0.085
Another method to measure the isotropy is Principle Components (IPC) Mu et al. (2018). To address the numerical instability issues when using PCA for isotropy measurement in high-dimensional embedding spaces, we use an approximation method for computing the IPC.
Given a normalized embedding matrix $E\in\mathbb{R}^{n\times d}$ where each row represents a unit vector, instead of computing the eigenvectors of the covariance matrix, we randomly sample $k$ unit vectors $\bm{u}_{1},\bm{u}_{2},...,\bm{u}_{k}$ from the unit hypersphere $\mathbb{S}^{d-1}$. For each sampled direction $\bm{u}_{i}$, we compute the function $F(\bm{u}_{i})=\sum_{j=1}^{n}\exp(\bm{e}_{j}{}^{\top}\bm{u}_{i})$, where $\bm{e}_{j}$ represents the $j$-th embedding vector. The approximated IPC is then defined as the ratio between the minimum and maximum values of $F$, given by (10).
$\mathrm{IPC}{\mathrm{approx}}=\frac{\min_{i\in{1,...,k}}F(u_{i})}{\max_{i\in{1,...,k}}F(u_{i})}$ (10)
If the approximated IPC value is close to 1, it means that the embedding space shows high isotropy where vectors are uniformly distributed across all directions in the high-dimensional space instead of clustering in some dominant directions. Conversely, when the IPC value approaches 0, it means significant anisotropy in the embedding space. We use $k=1000$ random directions to compute the approximation of IPC. Table 5 shows that the post-processing step of subtracting the condition c gives higher approximated IPC values, indicating the improvement of isotropy.
Model IPC (-c) IPC
NV 0.9611 0.9471
SFR 0.9463 0.9266
GTE 0.9490 0.9310
E5 0.8906 0.8368
SimCSE_base 0.9461 0.9297
SimCSE_large 0.9499 0.9406
Table 5: Approximated IPC values for each model. The left column of IPC (-c) lists values with post-processing step of subtracting the condition c.
Appendix C Full Results for all Models
Each LLM encoder has its own preferred pooling method, recommended by the original authors of those models. From Table 6 we see that the performance varies significantly depending on the pooling method being used, while the latent attention pooling used in NV reporting the best results. Note that NV does not support last or average pooling, while latent pooling is not supported by the other models.
Model Pooling sent/cond? Spear. Acc.
NV latent sent - c 16.98 52.24
latent sent 22.07 59.89
latent cond - c 31.32 64.91
latent cond 27.02 48.02
SFR last sent - c 11.88 46.97
last sent -0.18 33.77
last cond - c 19.28 50.66
last cond 13.00 44.33
average sent - c 19.54 57.52
average sent 11.89 43.01
average cond - c 20.38 52.51
average cond 18.32 46.44
average sent - c 19.54 57.52
average sent 11.89 43.01
average cond - c 20.38 52.51
average cond 18.32 46.44
GTE last sent - c 7.16 42.48
last sent 7.16 37.20
last cond - c 20.40 54.08
last cond 16.58 45.12
average sent - c 13.01 42.48
average sent 11.34 36.67
average cond - c 17.87 42.22
average cond 18.42 42.22
E5 average sent - c 11.08 42.21
average sent 3.77 33.77
average cond - c 12.01 42.74
average cond 6.18 37.47
last sent - c 9.28 41.69
last sent 0.49 31.13
last cond - c 8.90 39.84
last cond 3.05 34.56
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59 37.46
$\operatorname{CONC}(c+s)$ 4.00 35.09
$\operatorname{CONC}(s+c)$ c 8.32 34.56
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ c 8.58 48.02
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37 39.81
$\operatorname{CONC}(c+s)$ 1.25 34.82
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ c 7.05 42.74
$\operatorname{CONC}(c+s)$ c 6.00 36.67
Table 6: Spearman and accuracy scores for sentence embedding models. The original dimensionality of each model is indicated in parentheses. latent denotes latent attention pooling for NV, whereas last and average correspond to last token pooling and average pooling, respectively.
Appendix D Comparisons against Prior work
In Table 7, we compare our proposed method (CASE) against the best performances reported by the previously published C-STS methods. We report the results directly from their original publications and do not reproduce those methods in our comparisons. Moreover, when there are multiple encoder settings (e.g. bi-encoder, tri-encoder or cross-encoder) considered in prior work, we report the result for the best setting. The methods shown in the first four rows in Table 7 (Deshpande et al. (2023), Liu et al. (2025), Li et al. (2024b) and Yoo et al. (2024)) use the same training, validation and test sets from Deshpande et al. (2023). The penultimate method (Tu et al. (2024)) re-annotates the validation dataset from Deshpande et al. (2023) and further split it into a train/test (70% vs. 30%). They then use this train split for training their method and evaluate the performance on their test split.
Our proposed CASE method uses the same re-annotated dataset by Tu et al. (2024), but with a different random split of 70-30 for train and test purposes because Tu et al. (2024) did not released their train/test random split. Due to the differences in encoder settings and train/test data splits used in prior work, direct comparisons of Spearman correlation coefficients is difficult. However, despite these discrepancies, we see that Tu et al. (2024) and CASE report significantly higher Spearman correlations compared to the rest of the methods.
Method Spearman
Deshpande et al. (2023) (bi-encoder) 47.5
Liu et al. (2025) (bi-encoder) 48.1
Li et al. (2024b) (cross-encoder) 43.8
Yoo et al. (2024) (tri-encoder) 39.6
Tu et al. (2024) (bi-encoder) 75.9
Our CASE (bi-encoder) 70.0
Table 7: Spearman correlation of previously proposed methods and ours (CASE). We report the best performing encoder setting (shown in brackets) from each published paper.
Appendix E Additional Qualitative Analysis
cos(s1, s2)   $0.5006\rightarrow 0.9016$
Condition 1: The color of the dress. Condition 2: The name of the game.
Answer 1: orangeAnswer 2: grayRating: 1 Answer 1: tennisAnswer 2: tennisRating: 5
cos(s1, s2; c1)   $0.4757\rightarrow 0.2522$ cos(s1, s2; c2)   $0.6233\rightarrow 0.9660$
cos(s1, orange; c1)   $0.1986\rightarrow 0.3551$ cos(s1, tennis; c2)   $0.1135\rightarrow 0.6448$
cos(s2, gray; c1)   $0.0559\rightarrow 0.6061$ cos(s2, tennis; c2)   $0.0983\rightarrow 0.6426$
s1: Two snow skiers with ski poles and snow skis, standing on top of a snow covered mountain with other skiers around them.
cos(s1, s2)   $0.6318\rightarrow 0.7702$
Condition 1: The number of person. Condition 2: The type of job.
Answer 1: twoAnswer 2: oneRating: 1 Answer 1: skierAnswer 2: skierRating: 5
cos(s1, s2; c1)   $0.6358\rightarrow 0.2788$ cos(s1, s2; c2)   $0.7502\rightarrow 0.9539$
cos(s1, two; c1)   $-0.0970\rightarrow 0.7014$ cos(s1, skier; c2)   $0.2488\rightarrow 0.8016$
cos(s2, one; c1)   $0.0227\rightarrow 0.7953$ cos(s2, skier; c2)   $0.3409\rightarrow 0.8032$
s1: A bunch of people standing around at the beach with a kite in the air.
cos(s1, s2)   $0.3988\rightarrow 0.5350$
Condition 1: The type of hobby. Condition 2: The type of location.
Answer 1: kite flyingAnswer 2: surfRating: 1 Answer 1: beachAnswer 2: beachRating: 5
cos(s1, s2; c1)   $0.4386\rightarrow 0.4886$ cos(s1, s2; c2)   $0.4825\rightarrow 0.8582$
cos(s1, kite flying; c1)   $0.3457\rightarrow 0.7717$ cos(s1, beach; c2)   $0.2254\rightarrow 0.7281$
cos(s2, surf; c1)   $0.1034\rightarrow 0.6665$ cos(s2, beach; c2)   $0.1129\rightarrow 0.6703$
Table 8: Example of similarity scores for two conditions applied to the same sentence pair, based on linear MLP with dimensionality $512$ on NV. The table shows how supervised MLP projection improves the CASE for C-STS task. cos(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes cosine similarity. Answer 1 and Answer 2 refer to the corresponding answers for Sentence 1 and Sentence 2 under conditions, respectively. The predicted similarity scores before and after applying supervised MLP are listed on the left and right of the arrow. That is, the similarity score for original high-dimensional LLM-based embeddings is on the left of the arrow, while the similarity score for CASE is on the right.
We briefly show an example of our method in Table 8. We treat the conditions as questions for sentences and extract the relevant information as answers to compare whether CASE can focus on the condition-related information. The overall similarity trend is consistent with the actual ratings. For all sentence and answer pairs, the similarity scores increase after supervised projection. This demonstrates that the embeddings effectively capture the conditional meaning shift with higher condition-dependent similarity.

### v4

@@ Acknowledgements
Our evaluations cover only English, which is a morphologically limited language. To the best of our knowledge, C-STS datasets have not been annotated for languages other than English, which has forced all prior work on C-STS to conduct experiments using only English data. However, the sentence encoders we used in our experiments support multiple languages. Therefore, we consider it to be an important future research direction to create multilingual datasets for C-STS and evaluate the effectiveness of our proposed method in multilingual settings.
There is a large number of sentence encoders (over 18,000 sentence encoders as in January 2026 evaluated in Hugging Face Hub [^4] https://huggingface.co/models). However, due to computational limitations, we had to select a subset covering the best performing (top-ranked on MTEB leaderboard at the time of writing) models for our evaluations.
@@
We did not collect or annotate any datasets in this project. Instead, we use existing C-STS datasets annotated and made available by Deshpande et al. (2023), Tu et al. (2024), and Zhang et al. (2025). To the best of our knowledge, no ethical issues have been raised regarding those datasets.
@@ Appendix A Models
Appendix B Ablation Study for Supervised FFN
Table 7 provides the ablation study of linear and non-linear FFN, using NV as a representative model (for other encoders, we observe similar trends). We assess the impact of including a dropout layer. For non-linear FFN, we further compare the effects of different activation functions (ReLU, GeLU, SiLU, LeakyReLU) and the number of layers (one, two). Note that based on preliminary experiments, we set the dropout rate to 20% for the Linear FFN and 15% for the Non-linear FFNs as the best configuration. Additionally, for two-layer non-linear FFNs, the intermediate layer dimension is set to 1024, which performs best among the candidate set {256, 768, 1024, 2048, 8192}.
Linear FFN without dropout gives higher performance than that with dropout, with a Spearman correlation of 73.55. For non-linear FFNs, one-layer FFNs generally outperform two-layer FFNs by 1-3 points. Additionally, introducing a dropout layer improves the performance, probably by reducing the impact of overfitting. Among the four activation functions, LeakyReLU yields the best results for both two-layer and one-layer FFNs, achieving a Spearman correlation over 74. Consequently, we proceed with the linear FFN (without dropout) and the non-linear FFN (one layer, activation function LeakyReLU, dropout of 15%) for the subsequent experiments.
Activation w/ Dropout w/o Dropout
[…]
72.90 73.55
[…]
   ReLU 73.93 71.47
   GeLU 67.66 69.06
   SiLU 69.33 69.58
   LeakyReLU 74.05 73.29
[…]
   ReLU 73.27 71.56
   GeLU 71.05 70.56
   SiLU 71.08 71.13
   LeakyReLU 74.94 71.16
Table 7: Ablation study comparing Linear and Non-linear FFNs. Numbers in parentheses indicate layer dimensions.
Appendix C Isotropy of Embeddings
[…]
@@ Appendix C Isotropy of Embeddings
From Table 8, Figure 5, and Figure 6, we see that the embeddings after subtracting $\bm{c}$ have a lower mean cosine similarity to the mean vector and a higher standard deviation, indicating that they are more spread out in the embedding space. In contrast, the embeddings without subtracting $\bm{c}$ are more clustered around a central direction (higher mean, lower standard deviation), reflecting anisotropy, a tendency for vectors to concentrate in a narrow region. Therefore, embeddings after subtracting c tend to be more isotropic, indicating better distributional diversity.
Model Embedding Type Mean Std
NV cond - c 0.407 0.084
cond 0.492 0.067
SFR cond - c 0.537 0.055
cond 0.708 0.021
GTE cond - c 0.489 0.067
cond 0.696 0.036
E5 cond - c 0.542 0.056
cond 0.897 0.010
SimCSE_base $\operatorname{CONC}(c+s)$ - c 0.254 0.095
$\operatorname{CONC}(c+s)$ 0.347 0.088
SimCSE_large $\operatorname{CONC}(c+s)$ - c 0.248 0.110
$\operatorname{CONC}(c+s)$ 0.333 0.085
We use an efficient approximation method to compute the isotropy Durdy et al. (2023). Given a normalized embedding matrix $E\in\mathbb{R}^{n\times d}$ where each row represents a unit vector, instead of computing the eigenvectors of the covariance matrix, we randomly sample $k$ unit vectors $\bm{u}_{1},\bm{u}_{2},...,\bm{u}_{k}$ from the unit hypersphere $\mathbb{S}^{d-1}$. For each sampled direction $\bm{u}_{i}$, we compute the function $F(\bm{u}_{i})=\sum_{j=1}^{n}\exp(\bm{e}_{j}{}^{\top}\bm{u}_{i})$, where $\bm{e}_{j}$ represents the $j$-th embedding vector. The estimated isotropy is then defined as the ratio between the minimum and maximum values of $F$, given by (5).
$I_{\text{iso}}\approx\frac{\min_{i\in{1,...,k}}F(u_{i})}{\max_{i\in{1,...,k}}F(u_{i})}$ (5)
$I_{\text{iso}}$ values close to 1 indicates that the embedding space shows high isotropy where vectors are uniformly distributed in all directions in the high-dimensional space instead of clustering in some small number of dominant directions. In contrast, when $I_{\text{iso}}$ approaches 0, it means a significant anisotropy in the embedding space. We use $k=1000$ random directions to compute the approximation of isotropy. Table 9 shows that the post-processing step of subtracting the condition c gives higher $I_{\text{iso}}$ values, indicating the improvement of isotropy. Intuitively, when the embedding space is isotropic it becomes easier to differentiate between smaller similarity differences, thus improving the C-STS estimation.
Model $I_{\text{iso}}$ (-c) $I_{\text{iso}}$
NV 0.9611 0.9471
SFR 0.9463 0.9266
GTE 0.9490 0.9310
E5 0.8906 0.8368
SimCSE_base 0.9461 0.9297
SimCSE_large 0.9499 0.9406
Table 9: Isotropy values $I_{\text{iso}}$ for each model. The left column of $I_{\text{iso}}$ (-c) lists values with post-processing step of subtracting the condition c.
Model Pooling sent/cond? Spear.
NV latent sent - c 16.98
latent sent 22.07
latent cond - c 37.61
latent cond 33.85
SFR last sent - c 11.88
last sent -0.18
last cond - c 19.28
last cond 13.00
average sent - c 19.54
average sent 11.89
average cond - c 20.44
average cond 18.32
GTE last sent - c 7.16
last sent 7.16
last cond - c 20.40
last cond 16.58
average sent - c 13.01
average sent 11.34
average cond - c 17.87
average cond 18.42
E5 average sent - c 11.08
average sent 3.77
average cond - c 15.37
average cond 6.07
last sent - c 9.28
last sent 0.49
last cond - c 8.90
last cond 3.05
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59
$\operatorname{CONC}(c+s)$ 4.00
$\operatorname{CONC}(s+c)$ - c 8.32
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ - c 8.58
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37
$\operatorname{CONC}(c+s)$ 1.25
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ - c 7.05
$\operatorname{CONC}(c+s)$ - c 6.00
Table 10: Spearman scores for sentence embedding models. The original dimensionality of each model is indicated in parentheses. latent denotes latent attention pooling for NV, whereas last and average correspond to last token pooling and average pooling, respectively.
Appendix D Full Results for all Models
Each LLM encoder has its own optimal pooling method, recommended by the original authors of those models. From Table 10 we see that the performance varies significantly depending on the pooling method being used, while the latent attention pooling used in NV reporting the best results. Note that NV does not support last or average pooling, while latent pooling is not supported by the other models.
@@ Appendix D Full Results for all Models
(c) SimCSE_base: $\operatorname{CONC}(c+s)$ - c
@@
(e) SimCSE_large: $\operatorname{CONC}(c+s)$ - c